OpenStudy (anonymous):

please answer: 1. Find the angle between the given vectors to the nearest tenth of a degree. u = <8, 4>, v = <9, -9> answers; 81.6° 25.8° 35.8° 71.6° 2. A building has a ramp to its front doors to accommodate the handicapped. If the distance from the building to the end of the ramp is 25 feet and the height from the ground to the front doors is 7 feet, how long is the ramp? (Round to the nearest tenth.) answers: 24.0 ft 9.9 ft 26.0 ft 5.7 ft 3. Solve the triangle. A = 52°, b = 10, c = 7

OpenStudy (ybarrap):

1. $$ u = <8, 4>, v = <9, -9> \\ u\cdot v=|u||v|\cos\theta\\ \theta=\arccos \cfrac{u\cdot v}{|u||v|} $$ Make sense?

OpenStudy (anonymous):

a little bit..

OpenStudy (ybarrap):

2.  $$ \large x=\sqrt{25^2-7^2} $$ Where x is length of ramp. Back to 1. $$ u\cdot v=u = <8, 4>\cdot<9, -9> =72-36=36\\ |u|=\sqrt{8^2+4^2}\\ |v|=\sqrt{9^2+(-9)^2}\\ $$

OpenStudy (ybarrap):

Is this the problem for 3? 3. Solve the triangle. A = 52°, b = 10, c = 7 

OpenStudy (anonymous):

so number 2 is A?

OpenStudy (anonymous):

2. Pythagorean theorem C. 26.0 ft